Project

General

Profile

Update 22 May 2017

IPAC '17 results; LH2 commissioning and MICE status
Added by Kyberd, Paul 2 months ago

New results presented at IPAC17

The MICE expedition to IPAC included contributions from Chris Rogers, Paulo Franchini, Chris Hunt and Tanaz Mohayai:
• Chris R presented the first results of our study of the evolution of single-particle amplitudes (http://micewww.pp.rl.ac.uk/projects/analysis/wiki/Publications_and_Figures_Step_IV#Study-of-Ionization-Cooling-with-the-MICE-Experiment-Preliminary). The goal now is to bring the analysis to the state where we can publish a “rapid communication” by the end of the summer;
• Paulo presented the Step IV emittance-measurement and scattering results. His poster contained a first pair of plots from the energy loss analysis (S. Wylbur, see http://micewww.pp.rl.ac.uk/projects/analysis/wiki/Publications_and_Figures_Step_IV#Energy-Loss-in-Lithium-Hydride-and-Hydrogen-Absorbers-Preliminary);
• Chris H presented the performance of the cooling demonstration configuration accepted for publication in PRAB and the simplified scenario developed as an upgrade to Step IV (see http://micewww.pp.rl.ac.uk/projects/analysis/wiki/Publications_and_Figures_Step_IV#Layout-of-the-MICE-Demonstration-of-Muon-Ionization-Cooling-Including-Simplified-Design); and
• Tanaz presented the MC study of her “kernel-density estimator” analysis.

Progress on the commissioning of the liquid-hydrogen system in the Hall

The commissioning programme for the liquid-hydrogen delivery system in the MICE Hall has been progressing steadily. The system was cooled down to operating temperature last week. To prove the system, the commissioning programme includes the liquefaction of neon using the full LH2 system. This test is now underway. Neon was introduced into the system on Friday. Measurements of system parameters indicate that the cooling power is more than sufficient but the power that can be supplied by the heaters on the second stage of the cryocooler is too low to allow the system to be regulated. The heater power can be increased by exchanging the voltage source that drives current through the heaters. This will be done tomorrow (Monday 22May17).

The immediate programme for the LH2-system commissioning is the verification of the control sequences. This will be carried out between Monday 22May17 and Wednesday 24May17. On Wednesday, the system will be warmed up so that the absorber can be moved to R9 and to allow the beam-line to be configured for data taking.

On the 16th and 17th May 2017, a full HAZOP of the system was carried out in the light of changes made in response to experience gained in operating the LH2 system in R9 and the operation of the superconducting-magnet channel. The HAZOP recommended a number of actions be completed before the system is operated with hydrogen. The necessary modifications will be implemented and the system reinstalled such that commissioning can be completed in June 2017.

Status of the MICE apparatus

The expert-led startup of the experiment on the 08May went well; with the exception of the tracker, all detectors and the beam line operated without major issues.

The tracker cryostats have been a cause of concern. It was decided to carry out routine maintenance on two of the four cold-heads. When “cryo-2” was cooled down after maintenance, a large number of new dead channels were observed in the readout. This is now believed to be due to the freezing of water. While the maintenance operations were underway, the cryocooler on cryo-4 failed and it was decided to warm up the remaining two cryostats to allow maintenance on their cold-heads to be carried out.

In the light of the number of bad channels generated in the cool-down of cryo-2, a careful study of the content of the gas in the insulating vacuum space was carried out. This work culminated when Alan Bross, Ed Overton and Craig Macwaters discovered substantial amounts of water vapour using a custom liquid-nitrogen cold trap that Alan brought to RAL from FNAL. Using the cold trap, a programme to remove the water vapour has been carried out successfully on cryo-3. The programme to “dry out” the remaining cryostats is now nearing completion and the tracker group anticipates that the tracker will be ready for data taking by the end of the week.

The MICE Operations Managers for this Cycle, Victoria Blackmore and Ed Overton, are doing a good and patient job of responding with their usual professionalism and enthusiasm to the challenges that the hardware has been presenting them with!

Plans for data taking in Cycle 2017/01
The shift schedule for the present User Cycle (2017/01) was arranged to culminate in 24/7 running for the last shift block of the Cycle (next week). Given the outcome of the LH2-system HAZOP, the modifications required to the LH2 system and the fact that the tracker will be ready for operation at the end of the coming week, the run plan has been revisited to optimise the use of the data-taking time.

Analysis of the emittance-evolution data sets indicates that careful studies will be required to understand the transmission through the cooling cell. In particular, we need to study carefully the apertures that limit the transmission or cause energy-loss and scattering that may bias our measurements. After consultation and some discussion, the Analysis Group has requested empty-channel data in the magnet configurations in which the emittance-evolution data was taken. This data will allow us to study transmission in the absence of the effect of the LiH absorber.

In the light of the considerations outlined above, the programme for the rest of the present Cycle will be:
• Continued evaluation of the LH2 delivery system with emphasis on proving the controls sequences: Monday (22May17)—Wednesday (24May17);
• Extraction of LH2 absorber, re-insertion of the FC and establishment of the magnetic channel: Thursday (25May17)—Saturday (27May17);
• Magnet commissioning: Sunday (28May17)—Monday (29May17); and
• Empty-channel data taking (Monday 29May17)—Friday (02Jun17).

All hands meeting

Through the STFC MICE-UK Oversight Committee and communications from STFC to the MICE-UK PI (Paul Soler) the STFC has indicated that its financial position is such that, unless there is a change of circumstance, it is unable to fund the upgrade to Step IV by which we proposed to deliver the demonstration of ionization cooling.

In the light of this development, our ongoing discussions with IHEP Protvino on the possibility of executing a cooling demonstration in the U70 accelerator and the change to the run plan for the present User Cycle, I would like to call an “All Hands” meeting at 15:00 (GMT) on Thursday 25May17. We will use the familiar “video conference” arrangements for the meeting; Melissa Uchida will communicate the details shortly. By using the usual VC slot, I hope that many of you will be able to attend.

With best wishes …

Ken


Comments